# From Perception to Reasoning: Deep Thinking Empowers Multimodal Large Language Models

### v1

@@ 1 Introduction
With the support of vast data and powerful computational resources, Multimodal Large Language Models (MLLMs) have made significant advances in the understanding and generation of cross-modal content (e.g., text, images, videos) in recent years(kim2021vilt; li2021align; li2022blip; yu2022coca; wang2023image; chen2023x; li2023blip; wang2023caption; bai2025qwen2; zhang2023video; liu2023visual; lu2024deepseek; dong2023dreamllm; lin2024video; yin2023survey; caffagni2024revolution; zhang2024mm). They have been widely applied to tasks such as Image Captioning, Visual Question Answering (VQA)(antol2015vqa), and Video Captioning(venugopalan2015sequence). Despite their impressive performance in perception and generation, MLLMs still exhibit notable shortcomings when confronted with tasks involving complex reasoning(ghaffari2024exploring; malkinski2024reasoning; shiri2024empirical; imam2025can). Specifically, current MLLMs primarily rely on implicit reasoning, which entails making predictions based on statistical patterns in training data(bai2024survey; wang2025multimodal), and they lack explicit, interpretable intermediate reasoning steps. Consequently, they are limited in their capabilities for multi-step logical inference, causal reasoning, and compositional generalization(lu2022learn; li2025perception).
@@
To address similar issues in Large Language Models (LLMs), researchers have introduced the Chain-of-Thought (CoT) reasoning mechanism(wei2022chain; kojima2022large; wang2022self), aiming to enhance their logical reasoning abilities. The core idea of Chain-of-Thought is to decompose complex problems into a series of explicit intermediate reasoning steps, thereby simulating the human process of constructing a logical chain step by step. This approach has demonstrated notable advantages in tasks involving arithmetic, commonsense, and logical reasoning, while also improving the interpretability and transparency of model decision-making(huang2022towards; chu2024navigate; xia2025beyond). In the LLM domain, representative models such as OpenAI o1 and DeepSeek-R1(guo2025deepseek)have integrated Chain-of-Thought strategies to achieve significant breakthroughs in mathematical and logical reasoning tasks, effectively advancing the reasoning capabilities of LLMs.
@@
Inspired by this successful paradigm, researchers have recently begun exploring the extension of Chain-of-Thought reasoning to MLLMs, leading to the development of Chain-of-Thought-based Multimodal Large Language Models (CoT-MLLMs)(zhang2023multimodal; chen2024m3cot). This emerging paradigm aims to embed structured reasoning steps into the multimodal modeling framework, enabling models to perform higher-level cross-modal logical reasoning when processing various modalities such as vision and language. By doing so, CoT-MLLMs can capture deep semantic associations across modalities, thereby improving overall reasoning performance and interpretability(bi2025reasoning; lin2025mind; chen2025towards). Compared to LLMs that deal with a single text modality, enhancing the reasoning capabilities of MLLMs poses additional challenges, as they must simultaneously handle multimodal information. Thus, establishing effective Chains of Thought across modalities becomes a critical challenge. This involves not only aligning information across modalities but also constructing hierarchical reasoning structures to support deep cross-modal inference.
@@
To synthesize the current state of development in the field, several surveys(wang2025multimodal; bi2025reasoning; li2025perception) have emerged, offering researchers a comprehensive overview. Building upon prior work, this survey aims to provide a differentiated perspective with a greater emphasis on in-depth theoretical analysis. In contrast to existing works that focus on summarizing technical approaches, the core contribution of this survey lies in further analyzing and discussing the underlying mechanism by which MCoT enhances the reasoning capabilities of models, thereby answering the key question of ”why it is effective.” Furthermore, in the sections concerning the evaluation of, as well as the challenges and future directions, this survey provides a more systematic classification and summary. In summary, this review systematically organizes the latest research progress in the field of CoT-MLLMs, covering its core methods, evaluation benchmarks and metrics, typical application domains, as well as existing challenges and future directions. We hope that the insights and synthesis provided in this work can offer a structured reference and theoretical foundation for this emerging research direction of CoT-MLLMs, thereby promoting the sustained development of the field.
@@ 2.1 Multimodal Large Language Model
In recent years, MLLMs have achieved remarkable progress in cross-modal understanding and generation tasks, giving rise to a number of representative models. For example, Flamingo(alayrac2022flamingo) demonstrates excellent performance in zero-shot and few-shot tasks through a vision-language bridging module; BLIP-2(li2023blip) employs a two-stage training strategy to achieve efficient vision-language alignment; LLaVA(liu2023visual) enhances multimodal reasoning capabilities by effectively adapting visual encoders to language models. In addition, GPT-4V, as the multimodal version of GPT-4, exhibits strong integrated reasoning abilities in image understanding tasks.
@@
The fundamental architecture of MLLMs typically comprises three core components: modality encoders, a modality fusion module, and a large language model. The modality fusion module is a key component of MLLMs, responsible for semantic alignment and information integration across different modalities to facilitate cross-modal understanding and generation. Common approaches include the Q-Former mechanism, adopted by models such as BLIP-2(li2023blip), Video-LLaMA (zhang2023video), and InstructBLIP(dai2023instructblip), as well as linear projection layers or multi-layer perceptrons (MLPs) used in models like MiniGPT-4(zhu2023minigpt), LLaVA(liu2023visual), and PMC-VQA(zhang2023pmc). Through the coordinated operation of these three modules, MLLMs are capable of inter-modal information interaction, enabling cross-modal reasoning and content generation.
@@
Currently, most MLLMs perform well on basic perception tasks—such as object recognition and image captioning—but still exhibit notable deficiencies in complex reasoning tasks. For instance, Yang et al.(yang2025r1) observed that while Qwen2.5-VL(bai2025qwen2) possesses certain multimodal processing capabilities, it tends to make reasoning errors involving spatial relationships and logical judgments in region coloring tasks, ultimately leading to incorrect answers. Similarly, Toh et al.(toh2025jumping) found that the OpenAI o1 model underperforms humans on the PUZZLEVQA dataset, which involves simple visual abstract reasoning, and that GPT-4o shows significant weaknesses in perceptual and inductive reasoning. The primary reason lies in the fact that most current MLLMs rely heavily on statistical pattern prediction, lacking explicit reasoning mechanisms. This makes them prone to jumping steps, misattribution, and other logical errors, rendering it difficult to replicate the step-by-step deductive reasoning process characteristic of human cognition.
@@
On the other hand, the visual perception capabilities of current MLLMs also remain incomplete. In particular, for reasoning-intensive tasks, there is often a disconnect between perception and understanding(wang2024mr; jiang2024chatrex; chen2025bring; yan2025visuriddles). Multimodal reasoning requires the model to efficiently extract, integrate, and organize information from multiple input sources such as images and text. However, existing models still face evident bottlenecks in this process. In experiments, Ghosal et al.(ghosal2024language) found that when models are provided with more accurate and structured perceptual information, their performance on visual reasoning tasks improves significantly—further highlighting the inadequacies in current MLLMs’ mechanisms for coordinating perception and reasoning.
@@ 2.2 CoT and M-CoT
To address the two major challenges faced by current MLLMs in complex reasoning tasks—namely, the lack of explicit and structured reasoning mechanisms, and the disconnection between perception and reasoning—researchers have introduced the CoT reasoning mechanism as an effective solution. By incorporating explicit intermediate reasoning steps, CoT enables models to simulate the step-by-step cognitive reasoning process of humans, and has been proven to significantly improve the performance of large models on complex cognitive tasks(wei2022chain).
@@
where $r_{t}$ denotes the textual reasoning process at step $t$. This method alleviates issues such as step-skipping and vague reasoning by incrementally constructing a reasoning path, and improves the traceability of the reasoning chain. Based on this, researchers have further developed various extended forms. For example, Long et al.(long2023large) and Yao et al.(yao2023tree) proposed Tree-of-Thought (ToT), which enhances exploration and evaluation capabilities by constructing branched reasoning paths; Chen et al. proposed Chain-of-Symbolic, which transforms natural language reasoning steps into formal symbolic operations, significantly improving model performance in mathematical proof tasks.
To further address the problem of ”disconnection between perception and reasoning”, researchers have extended the CoT paradigm to the multimodal domain, proposing Multimodal Chain-of-Thought (M-CoT), enabling collaborative reasoning across heterogeneous modalities(zhang2023multimodal). Figure 2 shows a comparative example between the black-box reasoning of conventional multimodal large models and a naive M-CoT reasoning method. It can be seen that the main advantage of M-CoT lies in the introduction of an explicit reasoning process. Some more advanced M-CoT methods further integrate modality perception into the multi-step reasoning process to achieve better collaborative reasoning across multiple modalities. M-CoT not only inherits the advantages of traditional CoT in modeling intermediate states and structured reasoning, but also introduces a cross-modal interaction mechanism, enabling joint perception and collaborative reasoning across modalities such as vision and language. Its reasoning process can be divided into three stages: the intra-modal semantic evolution stage, which depends not only on its own historical states and original input, but also on the regulation and guidance from the previous multimodal fusion state; the construction stage of cross-modal CoT fusion representation, where the current fusion state is based on all modalities’ current semantic states and the previous fusion state, further achieving semantic alignment; and the task-oriented output stage, which generates the final output based on the final state of the fusion reasoning chain, ensuring the coherence of the reasoning path and the interpretability of the answer. The above process can be formally expressed as:
@@ 3 Method: How CoT Empowers CoT-MLLMs?
Table 1: Summary of the methods and models of M-CoT.
@@
HoT(yao2023thinking) 2023-8 UCAS Image,Text FT Science QA
MaTCR(liu2023matcr) 2023-10 ICT CAS Image,Text FT Multimodal Generation
DDCoT(zheng2023ddcot) 2023-10 ShanghaiTech Image,Text CoT prompting Visual QA
@@
Chain-of-Look(xi2023chain) 2023-10 SUNY Buffalo Video,Text CoT prompting Medical Image Analysis
GCoT(ma2024dolphins) 2023-12 UW-Madison Video,Text Autonomous Driving
CoCoT(zhang2024cocot) 2024-1 UR Image,Text CoT prompting Multi-image Understanding
KAM-CoT(mondal2024kam) 2024-1 Samsung Image,Text Knowledge Augmentation Science QA
VCoT(rose2023visual) 2024-1 UCSB Image,Text Multi-domain
BBA(zhao2024bba) 2024-2 Tencent Image,Text Multi-domain
$M$-Retrieval(liu2023retrieval) 2024-3 XMU Image,Text Knowledge Augmentation Math, Science QA
CoS(liu2024chain) 2024-3 THU Image,Text Image QA
@@
CCoT(mitra2024compositional) 2024-4 UC Berkeley Image,Text CoT prompting Multi-domain
@@
MI-CoT(anand2024mm) 2024-4 IIIT-Delhi Image,Text CoT prompting Physics Reasoning
Cantor(gao2024cantor) 2024-4 XMU Image,Text Math, Science QA
VoT(fei2024video) 2024-5 NUS Video,Text Self-refinement Visual Understanding
Multimodal-CoT(zhang2023multimodal) 2024-5 SJTU Image,Text Science QA
IoT(zhou2024image) 2024-5 Westlake Image,Text CoT prompting Visual Understanding
CoTDiffusion(ni2024generate) 2024-6 TJU Image,Text Embodied AI
MC-CoT(tan2024boosting) 2024-7 Westlake Image,Text Self-consistency Visual QA
DCoT(jia2025dcot) 2024-9 NEU Image,Text CoT prompting Visual QA
CoT-ST(du2024cot) 2024-9 HIT Speech,Text FT Speech Translation
ARES(byun2024ares) 2024-10 OSU Image,Text FT + RL Visual QA
R-CoT(deng2024r) 2024-10 HUST Image,Text Geometric Reasoning
SKETCHPAD(hu2024visual) 2024-11 UW Image,Text Math Reasoning
LLaVA-o1(xu2024llava) 2024-11 THU Image,Text FT Beam Search Multimodal Reasoning
VIC(zheng2024thinking) 2024-11 UPenn Image,Text Multimodal Reasoning
Insight-V(dong2025insight) 2024-11 NTU Image,Text FT + RL Multimodal Reasoning
@@
VisVm(wang2024scaling) 2024-12 Microsoft Image,Text Search Method Multimodal Reasoning
@@
AtomThink(xiang2024atomthink) 2024-12 SYSU Image,Text FT Search Method Math Reasoning
EMMA-X(sun2024emma) 2024-12 SUTD Image,Text Embodied AI
@@
AR-MCTS(dong2024progressive) 2024-12 RUC Image,Text Tree Search Multi-domain
Mulberry(yao2024mulberry) 2024-12 NTU Image,Text FT Tree Search Multi-task
@@
LlamaV-o1(thawakar2025llamav) 2025-1 MBZUAI Image,Text FT Beam Search Multi-task
MVoT(li2025imagine) 2025-1 Microsoft Image,Text FT Spatial Reasoning
CoR(yu2025chain) 2025-1 THU Image,Text Math Reasoning
DeepSeek-R1(guo2025deepseek) 2025-1 DeepSeek Image,Text FT + RL Multimodal Reasoning
SpatialCoT(liu2025spatialcot) 2025-1 Huawei Image,Text FT Spatial Reasoning
Virgo(du2025virgo) 2025-2 RUC Image,Text FT Multimodal Reasoning
AStar(wu2025boosting) 2025-2 THU Image,Text Tree Search Multimodal Reasoning
LLaVA-CoT(xu2024llava) 2025-2 THU Image,Text FT Beam Search Multimodal Reasoning
MedVLM-R1(pan2025medvlm) 2025-2 ICL Image,Text RL Medical Image Analysis
Visual-RFT(liu2025visual) 2025-3 SJTU Image,Text RL Visual Perception
Audio-Reasoner(xie2025audio) 2025-3 NTU Audio,Text FT Audio Reasoning
@@
R1-Omni(zhao2025r1) 2025-3 Alibaba Video,Audio,Text RL Emotion Recognition
Vision-R1(huang2025vision) 2025-3 ECNU Image,Text FT + RL Math Reasoning
DriveLMM-o1(ishaq2025drivelmm) 2025-3 MBZUAI Image,Text FT Autonomous Driving
EmbodiedVSR(zhang2025embodiedvsr) 2025-3 Beijing HRC Image,Text Embodied AI
ICoT(gao2025interleaved) 2025-3 Soochow Image,Text CoT prompting Visual Reasoning
R1-Onevision(yang2025r1) 2025-3 ZJU Image,Text FT + RL Multimodal Reasoning
Embodied-Reasoner(zhang2025embodied) 2025-3 ZJU Image,Text Embodied AI
@@
VisuoThink(wang2025visuothink) 2025-4 FDU Image,Text Tree Search Geometric, Spatial Reasoning

### v2

@@ 1 Introduction
With the support of vast data and powerful computational resources, Multimodal Large Language Models (MLLMs) have made significant advances in the understanding and generation of cross-modal content (e.g., text, images, videos) in recent years(Kim et al., 2021; Li et al., 2021; 2022; Yu et al., 2022; Wang et al., 2023b; Chen et al., 2023; Li et al., 2023a; Wang et al., 2023a; Bai et al., 2025; Zhang et al., 2023a; Liu et al., 2023b; Lu et al., 2024; Dong et al., 2023; Lin et al., 2024; Yin et al., 2023; Caffagni et al., 2024; Zhang et al., 2024b). They have been widely applied to tasks such as Image Captioning, Visual Question Answering (VQA)(Antol et al., 2015), and Video Captioning(Venugopalan et al., 2015). Despite their impressive performance in perception and generation, MLLMs still exhibit notable shortcomings when confronted with tasks involving complex reasoning(Ghaffari & Krishnaswamy, 2024; Małkiński et al., 2024; Shiri et al., 2024; Imam et al., 2025). Specifically, current MLLMs primarily rely on implicit reasoning, which entails making predictions based on statistical patterns in training data(Bai et al., 2024; Wang et al., 2025f), and they lack explicit, interpretable intermediate reasoning steps. Consequently, they are limited in their capabilities for multi-step logical inference, causal reasoning, and compositional generalization(Lu et al., 2022; Li et al., 2025c).
@@
To address similar issues in Large Language Models (LLMs), researchers have introduced the Chain-of-Thought (CoT) reasoning mechanism(Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022), aiming to enhance their logical reasoning abilities. The core idea of Chain-of-Thought is to decompose complex problems into a series of explicit intermediate reasoning steps, thereby simulating the human process of constructing a logical chain step by step. This approach has demonstrated notable advantages in tasks involving arithmetic, commonsense, and logical reasoning, while also improving the interpretability and transparency of model decision-making(Huang & Chang, 2022; Chu et al., 2024; Xia et al., 2025b). In the LLM domain, representative models such as OpenAI o1 and DeepSeek-R1(Guo et al., 2025)have integrated Chain-of-Thought strategies to achieve significant breakthroughs in mathematical and logical reasoning tasks, effectively advancing the reasoning capabilities of LLMs.
@@
Inspired by this successful paradigm, researchers have recently begun exploring the extension of Chain-of-Thought reasoning to MLLMs, leading to the development of Chain-of-Thought-based Multimodal Large Language Models (CoT-MLLMs)(Zhang et al., 2023c; Chen et al., 2024c). This emerging paradigm aims to embed structured reasoning steps into the multimodal modeling framework, enabling models to perform higher-level cross-modal logical reasoning when processing various modalities such as vision and language. By doing so, CoT-MLLMs can capture deep semantic associations across modalities, thereby improving overall reasoning performance and interpretability(Bi et al., 2025; Lin et al., 2025d; Chen et al., 2025a). Compared to LLMs that deal with a single text modality, enhancing the reasoning capabilities of MLLMs poses additional challenges, as they must simultaneously handle multimodal information. Thus, establishing effective Chains of Thought across modalities becomes a critical challenge. This involves not only aligning information across modalities but also constructing hierarchical reasoning structures to support deep cross-modal inference.
@@
To synthesize the current state of development in the field, several surveys(Wang et al., 2025f; Bi et al., 2025; Li et al., 2025c) have emerged, offering researchers a comprehensive overview. Building upon prior work, this survey aims to provide a differentiated perspective with a greater emphasis on in-depth theoretical analysis. In contrast to existing works that focus on summarizing technical approaches, the core contribution of this survey lies in further analyzing and discussing the underlying mechanism by which MCoT enhances the reasoning capabilities of models, thereby answering the key question of ”why it is effective.” Furthermore, in the sections concerning the evaluation of, as well as the challenges and future directions, this survey provides a more systematic classification and summary. In summary, this review systematically organizes the latest research progress in the field of CoT-MLLMs, covering its core methods, evaluation benchmarks and metrics, typical application domains, as well as existing challenges and future directions. We hope that the insights and synthesis provided in this work can offer a structured reference and theoretical foundation for this emerging research direction of CoT-MLLMs, thereby promoting the sustained development of the field.
@@ 2.1 Multimodal Large Language Model
In recent years, MLLMs have achieved remarkable progress in cross-modal understanding and generation tasks, giving rise to a number of representative models. For example, Flamingo(Alayrac et al., 2022) demonstrates excellent performance in zero-shot and few-shot tasks through a vision-language bridging module; BLIP-2(Li et al., 2023a) employs a two-stage training strategy to achieve efficient vision-language alignment; LLaVA(Liu et al., 2023b) enhances multimodal reasoning capabilities by effectively adapting visual encoders to language models. In addition, GPT-4V, as the multimodal version of GPT-4, exhibits strong integrated reasoning abilities in image understanding tasks.
@@
The fundamental architecture of MLLMs typically comprises three core components: modality encoders, a modality fusion module, and a large language model. The modality fusion module is a key component of MLLMs, responsible for semantic alignment and information integration across different modalities to facilitate cross-modal understanding and generation. Common approaches include the Q-Former mechanism, adopted by models such as BLIP-2(Li et al., 2023a), Video-LLaMA (Zhang et al., 2023a), and InstructBLIP(Dai et al., 2023), as well as linear projection layers or multi-layer perceptrons (MLPs) used in models like MiniGPT-4(Zhu et al., 2023), LLaVA(Liu et al., 2023b), and PMC-VQA(Zhang et al., 2023b). Through the coordinated operation of these three modules, MLLMs are capable of inter-modal information interaction, enabling cross-modal reasoning and content generation.
@@
Currently, most MLLMs perform well on basic perception tasks—such as object recognition and image captioning—but still exhibit notable deficiencies in complex reasoning tasks. For instance, Yang et al.(Yang et al., 2025c) observed that while Qwen2.5-VL(Bai et al., 2025) possesses certain multimodal processing capabilities, it tends to make reasoning errors involving spatial relationships and logical judgments in region coloring tasks, ultimately leading to incorrect answers. Similarly, Toh et al.(Toh et al., 2025) found that the OpenAI o1 model underperforms humans on the PUZZLEVQA dataset, which involves simple visual abstract reasoning, and that GPT-4o shows significant weaknesses in perceptual and inductive reasoning. The primary reason lies in the fact that most current MLLMs rely heavily on statistical pattern prediction, lacking explicit reasoning mechanisms. This makes them prone to jumping steps, misattribution, and other logical errors, rendering it difficult to replicate the step-by-step deductive reasoning process characteristic of human cognition.
@@
On the other hand, the visual perception capabilities of current MLLMs also remain incomplete. In particular, for reasoning-intensive tasks, there is often a disconnect between perception and understanding(Wang et al., 2024a; Jiang et al., 2024; Chen et al., 2025b; Yan et al., 2025a). Multimodal reasoning requires the model to efficiently extract, integrate, and organize information from multiple input sources such as images and text. However, existing models still face evident bottlenecks in this process. In experiments, Ghosal et al.(Ghosal et al., 2024) found that when models are provided with more accurate and structured perceptual information, their performance on visual reasoning tasks improves significantly—further highlighting the inadequacies in current MLLMs’ mechanisms for coordinating perception and reasoning.
@@ 2.2 CoT and M-CoT
To address the two major challenges faced by current MLLMs in complex reasoning tasks—namely, the lack of explicit and structured reasoning mechanisms, and the disconnection between perception and reasoning—researchers have introduced the CoT reasoning mechanism as an effective solution. By incorporating explicit intermediate reasoning steps, CoT enables models to simulate the step-by-step cognitive reasoning process of humans, and has been proven to significantly improve the performance of large models on complex cognitive tasks(Wei et al., 2022).
@@
where $r_{t}$ denotes the textual reasoning process at step $t$. This method alleviates issues such as step-skipping and vague reasoning by incrementally constructing a reasoning path, and improves the traceability of the reasoning chain. Based on this, researchers have further developed various extended forms. For example, Long et al.(Long, 2023) and Yao et al.(Yao et al., 2023b) proposed Tree-of-Thought (ToT), which enhances exploration and evaluation capabilities by constructing branched reasoning paths; Chen et al. proposed Chain-of-Symbolic, which transforms natural language reasoning steps into formal symbolic operations, significantly improving model performance in mathematical proof tasks.
To further address the problem of ”disconnection between perception and reasoning”, researchers have extended the CoT paradigm to the multimodal domain, proposing Multimodal Chain-of-Thought (M-CoT), enabling collaborative reasoning across heterogeneous modalities(Zhang et al., 2023c). Figure 2 shows a comparative example between the black-box reasoning of conventional multimodal large models and a naive M-CoT reasoning method. It can be seen that the main advantage of M-CoT lies in the introduction of an explicit reasoning process. Some more advanced M-CoT methods further integrate modality perception into the multi-step reasoning process to achieve better collaborative reasoning across multiple modalities. M-CoT not only inherits the advantages of traditional CoT in modeling intermediate states and structured reasoning, but also introduces a cross-modal interaction mechanism, enabling joint perception and collaborative reasoning across modalities such as vision and language. Its reasoning process can be divided into three stages: the intra-modal semantic evolution stage, which depends not only on its own historical states and original input, but also on the regulation and guidance from the previous multimodal fusion state; the construction stage of cross-modal CoT fusion representation, where the current fusion state is based on all modalities’ current semantic states and the previous fusion state, further achieving semantic alignment; and the task-oriented output stage, which generates the final output based on the final state of the fusion reasoning chain, ensuring the coherence of the reasoning path and the interpretability of the answer. The above process can be formally expressed as:
@@ 3 Method: How CoT Empowers CoT-MLLMs?
Table 1: Summary of the methods and models of MCoT.
@@
HoT(Yao et al., 2023a) 2023-8 UCAS Image,Text FT Science QA
MaTCR(Liu et al., 2023d) 2023-10 ICT CAS Image,Text FT Multimodal Generation
DDCoT(Zheng et al., 2023) 2023-10 ShanghaiTech Image,Text CoT prompting Visual QA
@@
Chain-of-Look(Xi et al., 2023) 2023-10 SUNY Buffalo Video,Text CoT prompting Medical Image Analysis
GCoT(Ma et al., 2024b) 2023-12 UW-Madison Video,Text Autonomous Driving
CoCoT(Zhang et al., 2024a) 2024-1 UR Image,Text CoT prompting Multi-image Understanding
KAM-CoT(Mondal et al., 2024) 2024-1 Samsung Image,Text Knowledge Augmentation Science QA
VCoT(Rose et al., 2023) 2024-1 UCSB Image,Text Multi-domain
BBA(Zhao et al., 2024a) 2024-2 Tencent Image,Text Multi-domain
MM-Retrieval(Liu et al., 2023a) 2024-3 XMU Image,Text Knowledge Augmentation Math, Science QA
CoS(Liu et al., 2024c) 2024-3 THU Image,Text Image QA
@@
CCoT(Mitra et al., 2024) 2024-4 UC Berkeley Image,Text CoT prompting Multi-domain
@@
MI-CoT(Anand et al., 2024b) 2024-4 IIIT-Delhi Image,Text CoT prompting Physics Reasoning
Cantor(Gao et al., 2024) 2024-4 XMU Image,Text Math, Science QA
VoT(Fei et al., 2024) 2024-5 NUS Video,Text Self-refinement Visual Understanding
Multimodal-CoT(Zhang et al., 2023c) 2024-5 SJTU Image,Text Science QA
IoT(Zhou et al., 2024a) 2024-5 Westlake Image,Text CoT prompting Visual Understanding
CoTDiffusion(Ni et al., 2024) 2024-6 TJU Image,Text Embodied AI
MC-CoT(Tan et al., 2024a) 2024-7 Westlake Image,Text Self-consistency Visual QA
DCoT(Jia et al., 2025b) 2024-9 NEU Image,Text CoT prompting Visual QA
CoT-ST(Du et al., 2024) 2024-9 HIT Speech,Text FT Speech Translation
ARES(Byun et al., 2024) 2024-10 OSU Image,Text FT + RL Visual QA
R-CoT(Deng et al., 2024) 2024-10 HUST Image,Text Geometric Reasoning
SKETCHPAD(Hu et al., 2024b) 2024-11 UW Image,Text Math Reasoning
LLaVA-o1(Xu et al., 2024a) 2024-11 THU Image,Text FT Beam Search Multimodal Reasoning
VIC(Zheng et al., 2024b) 2024-11 UPenn Image,Text Multimodal Reasoning
Insight-V(Dong et al., 2025) 2024-11 NTU Image,Text FT + RL Multimodal Reasoning
@@
VisVm(Wang et al., 2024g) 2024-12 Microsoft Image,Text Search Method Multimodal Reasoning
@@
AtomThink(Xiang et al., 2024a) 2024-12 SYSU Image,Text FT Search Method Math Reasoning
EMMA-X(Sun et al., 2024c) 2024-12 SUTD Image,Text Embodied AI
@@
AR-MCTS(Dong et al., 2024) 2024-12 RUC Image,Text Tree Search Multi-domain
Mulberry(Yao et al., 2024) 2024-12 NTU Image,Text FT Tree Search Multi-task
@@
LlamaV-o1(Thawakar et al., 2025) 2025-1 MBZUAI Image,Text FT Beam Search Multi-task
MVoT(Li et al., 2025a) 2025-1 Microsoft Image,Text FT Spatial Reasoning
CoR(Yu et al., 2025) 2025-1 THU Image,Text Math Reasoning
DeepSeek-R1(Guo et al., 2025) 2025-1 DeepSeek Image,Text FT + RL Multimodal Reasoning
SpatialCoT(Liu et al., 2025b) 2025-1 Huawei Image,Text FT Spatial Reasoning
Virgo(Du et al., 2025) 2025-2 RUC Image,Text FT Multimodal Reasoning
AStar(Wu et al., 2025) 2025-2 THU Image,Text Tree Search Multimodal Reasoning
LLaVA-CoT(Xu et al., 2024a) 2025-2 THU Image,Text FT Beam Search Multimodal Reasoning
MedVLM-R1(Pan et al., 2025) 2025-2 ICL Image,Text RL Medical Image Analysis
Visual-RFT(Liu et al., 2025e) 2025-3 SJTU Image,Text RL Visual Perception
Audio-Reasoner(Xie et al., 2025) 2025-3 NTU Audio,Text FT Audio Reasoning
@@
R1-Omni(Zhao et al., 2025b) 2025-3 Alibaba Video,Audio,Text RL Emotion Recognition
Vision-R1(Huang et al., 2025) 2025-3 ECNU Image,Text FT + RL Math Reasoning
DriveLMM-o1(Ishaq et al., 2025) 2025-3 MBZUAI Image,Text FT Autonomous Driving
EmbodiedVSR(Zhang et al., 2025f) 2025-3 Beijing HRC Image,Text Embodied AI
ICoT(Gao et al., 2025) 2025-3 Soochow Image,Text CoT prompting Visual Reasoning
R1-Onevision(Yang et al., 2025c) 2025-3 ZJU Image,Text FT + RL Multimodal Reasoning
Embodied-Reasoner(Zhang et al., 2025e) 2025-3 ZJU Image,Text Embodied AI
@@
VisuoThink(Wang et al., 2025h) 2025-4 FDU Image,Text Tree Search Geometric, Spatial Reasoning